OpenStudy (anonymous):

Find the equation in standard form, of a line passing through (3, -4) and (5,1). a. 5x+2y=23 b. y=5/2x-23/2 c. 5x-2y=23 d. 2y-3y=9 I feel like the answer is in between A and C .... I'm leaning on C.

OpenStudy (anonymous):

use the general equation of the straight line \[\frac{ y-y _{1} }{ x-x _{1} }=\frac{ y _{2} -y _{1}}{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

okay, I got, -5/2 as the slope. then, I used point-slope form and got the equation y-(-4)=-5/2(x-3) so, I now have y= -5/2x-7

OpenStudy (anonymous):

Standard form is Ax+By=C with no fractions. It looks like you have the slope correct, but you did not multiply the -5/2 but the -3 when you simplified your equation

OpenStudy (johnweldon1993):

Try the slope 1 more time... 1 - (-4) 1 + 4 5 ------- = ------- = ---- Positive slope not negative 5 - 3 5 - 3 2

OpenStudy (anonymous):

So it would be A?

OpenStudy (johnweldon1993):

Then by using the point (5,1) we have \[\large y - 1 = \frac{5}{2}(x - 5)\] \[\large y - 1 = \frac{5}{2}x - \frac{25}{2}\] \[\large y = \frac{5}{2}x - \frac{23}{2}\] This is slope intercept form...how would you make this standard form?

OpenStudy (johnweldon1993):

*Hint..No fractions...

OpenStudy (anonymous):

multiplying? the denominators? or finding a way to cancel out 2.

OpenStudy (johnweldon1993):

Well by multiplying each term by 2 you essentially cancel the 2... \[\large (2y = \frac{5}{\cancel{2}}x - \frac{23}{\cancel{2}})\times \cancel{2} \] \[\large 2y = 5x - 23\] Now what? Remember Standard form is in the form Ax + By = C